OpenStudy (anonymous):

25 • 5 • 5 • 4 500 2000 2500 5000

OpenStudy (a1234):

25*5 = 125 125*5 = 625 625*4 = ?

OpenStudy (anonymous):

25x5x5x4=2500 i think